OpenStudy (anonymous):

find the remainder of: x^5-x^4+x^3+2x^2-x+4 divided by x^3+1

OpenStudy (owlcoffee):

Use the theorem of the remainder and you'll obtain the remainder of the division without actually doing it.

OpenStudy (anonymous):

could you check my answer? i got x^2+3 to be the remainder

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (owlcoffee):

hmm, I got x^2-1 as a remainder, try checking if you didin't make any operatory mistakes.

OpenStudy (anonymous):

ill type up what i did one sec

OpenStudy (anonymous):

http://gyazo.com/fe443b7183b2571e9403d002a40c1a4a

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \normalsize \text{thats right , remainder will be}\hspace{.33em}\\~\\ \dfrac{x^2+3}{x^3+1}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

wait its all over x^3+1 ?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \normalsize \text{yes it will be}\hspace{.33em}\\~\\ \Huge \dfrac{x^2+3}{x^3+1}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

could you explain where you got the x^3+1, i only got my answer to be x^2+3

OpenStudy (anonymous):

waiiit nevermindd i understand

OpenStudy (anonymous):

its what we dividing by lol

OpenStudy (mathmath333):

dividend =quotient * divisor + remainder by this u can understand

OpenStudy (anonymous):

yea thank you :D :D